OpenStudy (laronjaxon):

The math team sold 42 hamburger meals and 15 ribs meals at the pep rally for a total of $243. At the game that night they sold 36 hamburger meals and 30 ribs meals for a total of $294. What is the cost of 1 rib meal?

OpenStudy (laronjaxon):

$1 $4 $5 $3

OpenStudy (shane_b):

Write out the two equations and solve for x and y: \[42x+15y=243\]\[36x+30y=294\]Do you know how to solve this?

OpenStudy (laronjaxon):

no :o

OpenStudy (shane_b):

Ok...do you at least understand how those 2 equations were created from the problem?

OpenStudy (laronjaxon):

yes i do :o

OpenStudy (shane_b):

Great. Now take one of them and solve it in terms of either x or y. Let's take the bottom one:\[x=\frac{294-30y}{36}\]With me so far?

OpenStudy (laronjaxon):

yess

OpenStudy (anonymous):

Let x = Hamburger meals Let y = Ribs meals \[42x+15y=243\] \[36x+30y=294\] Multiply the first equation by 2, thus obtaining: \[84x+30y=486\] \[36x+30y=294\] Subtract second equation from first equation, obtaining single equation: \[48x=192\] \[x=4\] Since we are looking for the value of y, substitute x into one of the equation: \[42(4)+15y=243\] \[168+15y=243\] \[15y=75\] \[y=5\] OR \[36(4)+30y=294\] \[144+30y=294\] \[30y=150\] \[y=5\] Cost of hamburger meal: $4 Cost of ribs meal: $5

OpenStudy (laronjaxon):

Thank u! :o!

OpenStudy (shane_b):

He did it a different way than I would have but the answer is still correct. If want I'll continue explaining my way but if not that's fine too :)